A man left 3/7 of his money to his wife,1/2 of the remainder to his brother and the rest of it to be divided equally among his three children.if each of his children received $ 400, how much did his wife receive?
Answer by mananth(16949)

let him leave x
wife 3/ 7 x
remainder 4/ 7 x
Brother = 1/ 2 * 4/7 x

Balance = 4/7 - 2/ 7 x
Balance = 2/ 7 x

2/ 7 x/ 3 = 400
2/21 x= 400
x= 4200

wife gets 3/7 * 4200
wife gets 1800